# Sina Baharlouei

### 06 Dec 2023 v1

@@ Abstract
Training and deploying machine learning models that meet fairness criteria for protected groups are fundamental in modern artificial intelligence. While numerous constraints and regularization terms have been proposed in the literature to promote fairness in machine learning tasks, most of these methods are not amenable to stochastic optimization due to the complex and nonlinear structure of constraints and regularizers. Here, the term “stochastic” refers to the ability of the algorithm to work with small mini-batches of data. Motivated by the limitation of existing literature, this paper presents a unified stochastic optimization framework for fair empirical risk minimization based on $f$-divergence measures ($f$-FERM). The proposed stochastic algorithm enjoys theoretical convergence guarantees. In addition, our experiments demonstrate the superiority of fairness-accuracy tradeoffs offered by $f$-FERM for almost all batch sizes (ranging from full-batch to batch size of one). Moreover, we show that our framework can be extended to the case where there is a distribution shift from training to the test data. Our extension is based on a distributionally robust optimization reformulation of $f$-FERM objective under $\ell_{p}$ norms as uncertainty sets. Again in this distributionally robust setting, $f$-FERM not only enjoys theoretical convergence guarantees but also outperforms other baselines in the literature in the tasks involving distribution shifts. An efficient stochastic implementation of $f$-FERM is publicly available  [^1] https://github.com/optimization-for-data-driven-science/f-FERM.
@@ 1 Introduction
Machine learning models are increasingly deployed in critical applications ranging from healthcare (Ahmad et al., 2018) to image processing (Krizhevsky et al., 2017), education to job recruitment (Boselli et al., 2018), and social networking to cybersecurity (Xin et al., 2018). Machine learning practitioners have adopted learning algorithms to fathom inherently difficult and crucial problems. However, naïve deployment of these models may lead to several serious shortcomings and ethical issues such as biased predictions against minority groups (Angwin et al., 2016; Buolamwini and Gebru, 2018), vulnerability to adversarial attacks (Madry et al., 2017; Carlini and Wagner, 2017; Baharlouei et al., 2023), and lack of generalizability to novel environments (Arjovsky et al., 2019). Consequently, it is of utmost importance to have reliable and trustworthy models that are, in particular, fair and comply with equality norms and provisions worldwide (Act, 1964; Elford, 2023).
@@
With the increasing concern for the trustworthiness of unchecked machine learning algorithms, a broad class of paradigms has been proposed to counteract and mitigate both the cause and effects of model unreliability. Imposing statistical independence between model output and particular input features is of interest in various domains, especially when the generalization of a trained model is based on a collection of spurious features present in the training dataset (Dwork et al., 2012; Hardt et al., 2016; Yan et al., 2017). These could be sensitive features like gender, race, age, and/or income in the context of fairness, or could be confounding factors like environmental artifacts in the context of image classification (Arjovsky et al., 2019). Existing literature on imposing statistical independence between selected input features and model outputs is directed into three approaches: pre-processing, post-processing, and in-processing methods.
@@
Pre-processing methods entail upstream changes made in datasets to mask sensitive features or reduce the dependency of output variables on sensitive features through transforming data in a stage before the training phase (Kamiran and Calders, 2012; Zemel et al., 2013; Ustun et al., 2019). Post-processing methods involve model-specific adjustments to the model’s output to ensure the independence of predictions and sensitive attributes (Hardt et al., 2016; Alghamdi et al., 2022). While pre-processing and post-processing methods do not affect the training procedure, they fail to exploit underlying training mechanisms for the best achievable accuracy-fairness tradeoffs. Unsurprisingly enough, optimizing accuracy and fairness jointly (in-processing) leads to better tradeoffs than sequentially optimizing fairness and accuracy in a pre-processing or post-processing fashion.
@@
In-processing methods overcome these shortcomings of pre-processing and post-processing approaches by adding fairness constraints or regularizers to the training objective, penalizing dependence between sensitive attributes and output variables. (Zafar et al., 2017) utilizes covariance as the measure of independence between the sensitive attributes and the predictions. While such a measure is amenable to stochastic updates, it fails to capture correlations beyond linear. Alternatively, several non-linear measures such as Rényi correlation (Baharlouei et al., 2020), $\chi^{2}$ divergence (Lowy et al., 2022), $L_{\infty}$ distance (Donini et al., 2018), and Maximum Mean Discrepancy (MMD) (Prost et al., 2019) are proposed in the literature to establish the independence of the predictors and sensitive attributes. In-processing techniques can be model-specific (Wan et al., 2021; Aghaei et al., 2019) or generalizable to different training algorithms (Baharlouei et al., 2020; Lowy et al., 2022).
@@
While in-processing techniques outperform pre-processing and post-processing approaches, they are not scalable to large datasets because of a lack of adaptability to Stochastic Approximation methods (Mary et al., 2019; Lowy et al., 2022). From an optimization perspective, all aforementioned examples consist of regularization terms in their objective functions where the gradient cannot be described as a linear combination of data points functions. As a result, applying stochastic gradient descent or other stochastic first-order methods on the objective functions of such problems might not converge, especially for small batch sizes.
@@
Motivated by this, (Lowy et al., 2022) proposes a provably convergent stochastic optimization framework for Exponential Rényi Mutual Information as the measure of independence. More recently Zhong and Tandon (2023) use $f$-divergences as regularization terms to establish the independence between sensitive attributes and predictions. They estimate the $f$-divergence regularizers offline through multi-layer neural networks to avoid the computational challenges of devising scalable stochastic methods for nonconvex min-max problems. Our approach, on the other hand, directly solves the variational formulation for both full-batch and stochastic settings with convergence guarantees to non-spurious solutions. In Section 2, using the variational representation of $f$-divergences, we present a convergent stochastic optimization framework for fair learning via $f$-divergences. (Lowy et al., 2022) is a special case of $f$-divergences where $f(t)=t^{2}-1$ ($\chi^{2}$ divergence). Aside from $\chi^{2}$, all other divergences listed in Table 1 are not introduced in the literature to the best of our knowledge.
@@
The need of having convergent algorithms for fair empirical risk minimization does not end with designing methods amenable to stochastic approximation. Detection and mitigation of biases against protected groups in the presence of distribution shifts have been extensively studied in recent years. Lechner et al. (2021) theoretically shows that learning fair representations (pre-processing) is nearly impossible for the popular notions of fairness, such as demographic parity in the presence of the distribution shift. Ding et al. (2021), on the other hand, experimentally demonstrates that applying post-processing fairness techniques (Hardt et al., 2016) to learn fair predictors of income concerning race, gender, and age fails to transfer from one US state (training domain) to another state. Overlooking distribution shifts can lead to catastrophic decisions threatening the well-being of human subjects when deploying a trained model in certain hospitals to other hospitals . The current literature for handling distribution shifts with in-processing methods relies on certain assumptions on the type of distribution shift (demographic shift (Fang et al., 2020; Du and Wu, 2021; Maity et al., 2021; Giguere et al., 2021), label shift (Dai and Brown, 2020), and/or covariate shift (Rezaei et al., 2021; Singh et al., 2021)) or explicit access to the causal graph (Mishler and Dalmasso, 2022; Schrouff et al., 2022) of predictors, sensitive attributes, and target variables. As a result, they face practical limitations and cannot cope with most real-world problems involving complex shifts that cannot be categorized in the ones assumed in their works.
@@
Alternatively, Taskesen et al. (2020) provides convex objective functions for imposing fairness on logistic regression using constraint optimization. Staib and Jegelka (2019) use MMD for defining uncertainty sets around training distribution, whereas Husain (2020) use Integral Probability Measure (IPM) to mitigate the distribution shift. The main limitation of these approaches is their reliance on the convexity of the underlying learning model and lack of scalability due to incompatibility with stochastic optimization algorithms. Wang et al. (2023) uses the Maximum Mean Discrepancy (MMD) distance between the spectral norm of the Hessian matrix at advantaged and disadvantaged data points. However, they do not provide convergence guarantees for their proposed algorithm to any notion of optimality. In addition, the method is not necessarily amenable to stochastic updates. While we naturally define the uncertainty set directly on the joint distribution of sensitive attributes and predictions, they use the curvature of the obtained solution quantified by the norm of the Hessian matrix as a heuristic for promoting the robustness of the fair solution. In Section 3, we present our approach for fair inference in the presence of the distribution shifts in detail.
@@
Paper Organization: We structure our response towards designing scalable, robust, and fair algorithms into two sections. Section 2 motivates the design of unbiased gradient estimators of objectives with information-theoretic $f$-divergence regularizers. In Section 3, we present our approach for fair inference in the presence of the distribution shift in detail. We consider additional cases when prior knowledge of bounds on distributional distance is present. Our experiments provide an extensive examination of various $f$-divergences and their suitability as regularizers and also show the consistency of our method across all batch sizes in contrast to existing benchmarks. Similar experiments are carried out for robust training on varying amounts of distributional shifts in data.
@@ 2 Fair Empirical Risk Minimization via $f$-divergences
$\min_{{\bm{\theta}}}\quad\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\mathcal{D}\Big{(}\mathbb{P}({\hat{y}_{{\bm{\theta}}}({\mathbf{x}}),s}),\mathbb{P}(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}))\otimes\mathbb{P}(s)\Big{)},$ (1)
where ${\bm{\theta}}$ is the parameter to be learned (e.g. weights of the neural network); ${\mathbf{x}}_{i}\in\mathbb{R}^{d}$ is the $i$-th input feature vector; $y_{i}$ is the actual label/class for sample $i$; $\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i})$ is the prediction of the model for sample $i$; and $\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})$ is the loss function measuring the “goodness-of-fit” for sample $i$. Here, $\mathcal{D}$ is a divergence between the joint probability distribution of the predictions and sensitive attributes and the Kronecker product of their marginal distributions. Recall that $\hat{y}_{{\bm{\theta}}}$ and $s$ are statistically independent iff $\mathbb{P}({\hat{y}_{{\bm{\theta}}}({\mathbf{x}}),s})$ follows $\mathbb{P}(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}))\otimes\mathbb{P}(s)$. Therefore, the second term in (1) is zero if and only if $\hat{y}_{{\bm{\theta}}}$ and $s$ are statistically independent (complete fairness under the demographic parity notion).
Various divergence measures have been used in the algorithmic fairness literature to establish the independence of sensitive attributes and predictions. (Zafar et al., 2017) adopts covariance of sensitive attributes and predictions decision boundary as the measure of independence. While this choice of regularizer leads to an almost efficient optimization procedure, unfortunately, it cannot capture non-linear correlations between sensitive attributes and predictions. To address this issue, (Mary et al., 2019; Baharlouei et al., 2020) utilize Rényi correlation as the measure of independence. While these papers derive convergent algorithms for full-batch gradient descent, they fail to guarantee convergence for stochastic gradient descent (small-batch). Therefore, they are not scalable to large datasets. (Lowy et al., 2022) addresses this issue by introducing the Exponential Rényi Mutual Information (ERMI) which is equivalent to $\chi^{2}$ divergence. It is known that $\chi^{2}$ divergence falls under the general class of $f$-divergence measures (Nielsen and Nock, 2013). Therefore, a natural question is whether their framework can be naturally generalized to the class of $f$-divergence measures; and if so, whether more efficient fairness-imposing regularizers can be discovered in this class.
This section studies the fair empirical risk minimization regularized by a broad class of $f$-divergence measures. Let $\mathbb{P}$ and $\mathbb{Q}$ be two discrete probability measures taking values in $\mathcal{P}=\{1,\dots,m\}$. The $f$-divergence between $\mathbb{P}$ and $\mathbb{Q}$ is defined as (Polyanskiy and Wu, 2022, Def 4.9)(see Appendix B for the general continuous case):
@@
The above definition covers many well-known divergence measures (used for imposing fairness), such as KL-divergence for the choice of $f(t)=t\log(t)$ (Shui et al., 2022), or $\chi^{2}$ divergence when $f(t)=(t-1)^{2}$ (Lowy et al., 2022). As shown in Appendix C, $\mathcal{D}_{f}$ in (1) is zero if and only if the probability distribution of $s$ and $\hat{y}_{{\bm{\theta}}}$ are statistically independent for the choices of $f$ listed in Table 1. In addition, we prove that these $f$-divergences either cover or provide upper bounds for the popular notions of fairness violations in the literature, such as $\ell_{p}$ distances, Rényi correlation (Baharlouei et al., 2020), and demographic parity (equalized odds) violation. This means that by minimizing these regularizers, we are minimizing an upper bound of (other) popular fairness violation measures, and thus we are controlling them implicitly. Further, unlike Rényi correlation (Baharlouei et al., 2020; Grari et al., 2020), we can utilize Legendre-Fenchel duality (and variational representation) to develop (provably) convergent algorithms with stochastic (mini-batch) updates. This formulation and the resulting stochastic optimization algorithm are described in the next subsection.

### 07 Apr 2024 v2

@@ Abstract
Training and deploying machine learning models that meet fairness criteria for protected groups are fundamental in modern artificial intelligence. While numerous constraints and regularization terms have been proposed in the literature to promote fairness in machine learning tasks, most of these approaches are not amenable to stochastic optimization due to the complex and nonlinear structure of constraints and regularizers. Here, the term “stochastic” refers to the ability of the algorithm to work with small mini-batches of data. Motivated by the limitation of existing literature, this paper presents a unified stochastic optimization framework for fair empirical risk minimization based on $f$-divergence measures ($f$-FERM). The proposed stochastic algorithm enjoys theoretical convergence guarantees. In addition, our experiments demonstrate the superiority of fairness-accuracy tradeoffs offered by $f$-FERM for almost all batch sizes (ranging from full-batch to batch size of one). Moreover, we show that our framework can be extended to the case where there is a distribution shift from training to the test data. Our extension is based on a distributionally robust optimization reformulation of $f$-FERM objective under $\ell_{p}$ norms as uncertainty sets. Again, in this distributionally robust setting, $f$-FERM not only enjoys theoretical convergence guarantees but also outperforms other baselines in the literature in the tasks involving distribution shifts. An efficient stochastic implementation of $f$-FERM is publicly available  [^1] https://github.com/optimization-for-data-driven-science/f-FERM.
@@ 1 Introduction
Machine learning models are increasingly deployed in critical applications ranging from healthcare (Ahmad et al., 2018) to image processing (Krizhevsky et al., 2017), education to job recruitment (Boselli et al., 2018), and social networking to cybersecurity (Xin et al., 2018). Machine learning practitioners have adopted learning algorithms to fathom inherently difficult and crucial problems. However, naïve deployment of these models may lead to serious shortcomings such as biased predictions against minority groups (Angwin et al., 2016; Buolamwini & Gebru, 2018), vulnerability to adversarial attacks (Madry et al., 2017; Carlini & Wagner, 2017; Baharlouei et al., 2023), or lack of generalizability (Arjovsky et al., 2019). Consequently, it is of utmost importance to have reliable and trustworthy models that are, in particular, fair and comply with equality norms and provisions worldwide (Act, 1964; Elford, 2023).
@@
With the increasing concern for the trustworthiness of unchecked machine learning algorithms, a broad class of paradigms has been proposed to counteract and mitigate both the cause and effects of model unreliability. Imposing statistical independence between model output and particular input features is of interest in various domains, especially when the generalization of a trained model is based on a collection of spurious features present in the training dataset (Dwork et al., 2012; Hardt et al., 2016; Yan et al., 2017). These could be sensitive features like gender, race, age, and/or income in the context of fairness or confounding factors like environmental artifacts in the context of image classification (Arjovsky et al., 2019). Existing literature on imposing statistical independence between selected input features and model outputs is directed into three approaches: pre-processing, post-processing, and in-processing methods.
@@
Pre-processing methods entail upstream changes made in datasets to mask sensitive features or reduce the dependency of output variables on sensitive features through transforming data in a stage before the training phase (Kamiran & Calders, 2012; Zemel et al., 2013; Ustun et al., 2019). Post-processing methods involve model-specific adjustments to the model’s output to ensure the independence of predictions and sensitive attributes (Hardt et al., 2016; Alghamdi et al., 2022). While pre-processing and post-processing methods do not affect the training procedure, they fail to exploit underlying training mechanisms for the best achievable accuracy-fairness tradeoffs. Unsurprisingly enough, optimizing accuracy and fairness jointly (in-processing) leads to better tradeoffs than sequentially optimizing fairness and accuracy in a pre-processing or post-processing fashion.
@@
In-processing methods alternatively add fairness constraints or regularizers, penalizing dependence between sensitive attributes and output variables. (Zafar et al., 2017) utilizes covariance as the measure of independence between the sensitive attributes and the predictions. While such a measure is amenable to stochastic updates, it fails to capture correlations beyond linear. Alternatively, several non-linear measures such as Rényi correlation (Baharlouei et al., 2020), $\chi^{2}$ divergence (Lowy et al., 2022), $L_{\infty}$ distance (Donini et al., 2018), and Maximum Mean Discrepancy (MMD) (Prost et al., 2019) are proposed in the literature to establish the independence of the predictors and sensitive attributes. In-processing techniques can be model-specific (Wan et al., 2021; Aghaei et al., 2019) or generalizable to different training algorithms (Baharlouei et al., 2020; Lowy et al., 2022).
@@
While in-processing techniques outperform pre-processing and post-processing approaches, they are not scalable to large datasets because of a lack of adaptability to stochastic optimization (Mary et al., 2019; Lowy et al., 2022). All aforementioned examples consist of regularization terms in their objective functions where the gradient cannot be described as a linear combination of data point functions. As a result, applying stochastic gradient descent or other stochastic first-order methods on the objective functions of such problems might not converge, especially for small batch sizes.
@@
Motivated by this, (Lowy et al., 2022) proposes a stochastic optimization framework for Exponential Rényi Mutual Information as the measure of independence. More recently Zhong & Tandon (2023) use $f$-divergences as regularization terms to establish the independence between sensitive attributes and predictions. They estimate the $f$-divergence regularizers offline through multi-layer neural networks to avoid the computational challenges of devising scalable stochastic methods for nonconvex min-max problems. Our approach, on the other hand, directly solves the variational formulation for both full-batch and stochastic settings with convergence guarantees to non-spurious solutions. In Section 2, using the variational representation of $f$-divergences, we present a convergent stochastic optimization framework for fair learning via $f$-divergences. (Lowy et al., 2022) is a special case of $f$-divergences where $f(t)=t^{2}-1$ ($\chi^{2}$ divergence). Aside from $\chi^{2}$, all other divergences listed in Table 1 are not introduced in the literature to the best of our knowledge.
@@
Designing convergent stochastic algorithms for fair empirical risk minimization can be further explored in scenarios involving changes in the data distribution from the source to the target domain. Detection and mitigation of biases against protected groups in the presence of distribution shifts have been extensively studied in recent years. Lechner et al. (2021) theoretically shows that learning fair representations (pre-processing) is nearly impossible for the popular notions of fairness, such as demographic parity in the presence of the distribution shift. Ding et al. (2021), on the other hand, experimentally demonstrates that applying post-processing fairness techniques (Hardt et al., 2016) to learn fair predictors of income concerning race, gender, and age fails to transfer from one US state (training domain) to another state. Overlooking distribution shifts can lead to catastrophic decisions threatening the well-being of human subjects when deploying a trained model in certain hospitals to other hospitals (Schrouff et al., 2022). The current literature for handling distribution shifts with in-processing methods relies on certain assumptions on the type of distribution shift (demographic shift (Fang et al., 2020; Du & Wu, 2021; Maity et al., 2021; Giguere et al., 2021), label shift (Dai & Brown, 2020), and/or covariate shift (Rezaei et al., 2021; Singh et al., 2021)) or explicit access to the causal graph (Mishler & Dalmasso, 2022; Schrouff et al., 2022) of predictors, sensitive attributes, and target variables. As a result, they face practical limitations and cannot cope with most real-world problems involving complex shifts that cannot be categorized in the ones assumed in their works.
@@
Alternatively, Taskesen et al. (2020) provides convex objective functions for imposing fairness on logistic regression using constraint optimization. Staib & Jegelka (2019) use MMD for defining uncertainty sets around training distribution, whereas Husain (2020) use Integral Probability Measure (IPM) to mitigate the distribution shift. The main limitation of these approaches is their reliance on the convexity of the underlying learning model and lack of scalability due to incompatibility with stochastic optimization algorithms. Wang et al. (2023) uses the Maximum Mean Discrepancy (MMD) distance between the spectral norm of the Hessian matrix at advantaged and disadvantaged data points. However, they do not provide convergence guarantees for their proposed algorithm to any notion of optimality. In addition, the method is not necessarily amenable to stochastic updates. While we naturally define the uncertainty set directly on the joint distribution of sensitive attributes and predictions, they use the curvature of the obtained solution quantified by the norm of the Hessian matrix as a heuristic for promoting the robustness of the fair solution.
@@
Paper Organization: We structure our response towards designing scalable, robust, and fair algorithms into two sections. Section 2 motivates the design of unbiased gradient estimators of objectives with information-theoretic $f$-divergence regularizers. In Section 3, we present our approach for fair inference in the presence of the distribution shift in detail. Our experiments provide an extensive examination of various $f$-divergences and their suitability as regularizers and also show the consistency of our method across all batch sizes in contrast to existing benchmarks. Similar experiments are carried out for robust training on varying amounts of distributional shifts in data.
@@ 2 Fair Empirical Risk Minimization via $f$-divergences
$\vspace{-1mm}\min_{{\bm{\theta}}}\quad\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\mathcal{D}\Big{(}\mathbb{P}({\hat{y}_{{\bm{\theta}}}({\mathbf{x}}),s}),\mathbb{P}(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}))\otimes\mathbb{P}(s)\Big{)},$ (1)
where ${\bm{\theta}}$ is the learning parameters (e.g., weights of the neural network); ${\mathbf{x}}_{i}\in\mathbb{R}^{d}$ is the $i$-th input feature vector; $y_{i}$ is the actual label/class for sample $i$; $\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i})$ is the prediction of the model for sample $i$; and $\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})$ is the loss function measuring the “goodness-of-fit" for sample $i$. Here, $\mathcal{D}$ is a divergence between the joint probability distribution of the predictions and sensitive attributes and the Kronecker product of their marginal distributions. Recall that $\hat{y}_{{\bm{\theta}}}$ and $s$ are statistically independent iff $\mathbb{P}({\hat{y}_{{\bm{\theta}}}({\mathbf{x}}),s})$ follows $\mathbb{P}(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}))\otimes\mathbb{P}(s)$. Thus, the second term in (1) is zero iff $\hat{y}_{{\bm{\theta}}}$ and $s$ are statistically independent (complete fairness under the demographic parity notion).
This section studies the fair empirical risk minimization regularized by a broad class of $f$-divergence measures. Let $\mathbb{P}$ and $\mathbb{Q}$ be two discrete probability measures taking values in $\mathcal{P}=\{1,\dots,m\}$. The $f$-divergence between $\mathbb{P}$ and $\mathbb{Q}$ is defined as (Polyanskiy & Wu, 2022, Def 4.9)(see Appendix B for the general continuous case):
@@
The above definition, which is also known as $f$-mutual information (Lu et al., 2023; Csiszár, 1967), covers many known divergence measures used for imposing fairness, such as KL-divergence for the choice of $f(t)=t\log(t)$ (Shui et al., 2022), or $\chi^{2}$ divergence when $f(t)=(t-1)^{2}$ (Lowy et al., 2022). As shown in Appendix C, $\mathcal{D}_{f}$ in (1) is zero if and only if the probability distribution of $s$ and $\hat{y}_{{\bm{\theta}}}$ are statistically independent for the choices of $f$ listed in Table 1. In addition, we prove that these $f$-divergences either cover or provide upper bounds for the popular notions of fairness violations in the literature, such as $\ell_{p}$ distances, Rényi correlation (Baharlouei et al., 2020), and demographic parity (equalized odds) violation. This means that by minimizing these regularizers, we are minimizing an upper bound of (other) popular fairness violation measures, and thus we are controlling them implicitly. Further, unlike Rényi correlation (Baharlouei et al., 2020; Grari et al., 2020), we can utilize Legendre-Fenchel duality (and variational representation) to develop (provably) convergent algorithms with stochastic (mini-batch) updates. This formulation and the resulting stochastic optimization algorithm are described in the next subsection.